State, giving reasons, whether the following statements are True or False:
(1)The electric field associated with a plane electromagnetic wave is given by
E = 50 cos (ωt – 10πx) zˆ Vm^–1
. Its wavelength is 0.5 m.
1
Expert's answer
2020-02-24T10:49:42-0500

Answer: False

Its is lectric field associated with a plane electromagnetic wave. But its wavelength is


"\frac{2\pi}{\lambda}=10\pi"

"\lambda=\frac{2}{10}=0.2\ m"